40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Oct 21 18:40:46 2014, manasa, Update, General, End laser fiber setup 
    Reply  Fri Oct 24 10:08:24 2014, manasa, Update, General, Y AUX laser - fiber coupled 
       Reply  Fri Oct 24 18:53:23 2014, rana, Update, General, Y AUX laser - fiber coupled 
       Reply  Mon Oct 27 12:17:46 2014, manasa, Update, General, Y AUX laser coupling telescope design AUXmode.zipAUXmode.png
          Reply  Wed Oct 29 11:45:11 2014, manasa, Update, General, Y AUX laser - fiber coupled (52%) 31.png
Message ID: 10638     Entry time: Fri Oct 24 10:08:24 2014     In reply to: 10629     Reply to this: 10639   10640
Author: manasa 
Type: Update 
Category: General 
Subject: Y AUX laser - fiber coupled 

The Y end aux laser light leaking after the doubling crystal has been coupled into the 70m long PM fiber. 

Input power = 250mW;  Output after 70m = 20mW

The poor efficiency is partially due to the ellipticity of the beam itself and partially due to the compromise I had to make using a single lens to couple the light into the fiber (given the limitations in space). But 20mW should be more than sufficient for a beat note setup.

Light propagates as follows after the doubling crystal:

Doubler ---> Harmonic Separator (45deg) ---> Lens (f=12.5cm) --> Steering mirror (Y1) --> Fiber collimator ( Thor labs CFC-2X-C) --> FIber end

I will update photos of the setup shortly.

I have left the 70m fiber in its spool sitting at the Y end and blocked the light before the last Y1 steering mirror in the above setup. So it should be safe.

Other:

Through the course of the work, I disabled the ETMY oplev and enabled it before closing the enclosure. I also reduced the AUX laser power and brought it back up after the work.
I did a check to see if the arms are locking in both IR and green and they did.
 

ELOG V3.1.3-